# On the Dynamic Regret of Following the Regularized Leader: Optimism with History Pruning

### v1

@@ A.4 Dual-perspective
$r^{\star}(\bm{y})\doteq\sup_{\bm{x}}\langle{\bm{x}},{\bm{y}}\rangle-r(\bm{x}).$
@@
$\bm{x}^{\text{RL}}_{t+1}=\arg\min_{x}\langle\bm{g}_{1:t},\bm{x}\rangle+r_{0:t}(\bm{x})=\nabla r_{0:t}^{\star}(-\bm{g}_{1:t}),$
@@
$\bm{x}_{t+1}=\arg\min_{x}\langle\bm{p}_{1:t},\bm{x}\rangle+r_{0:t}(\bm{x})=\nabla r_{0:t}^{\star}(-\bm{p}_{1:t})=\nabla r_{0:t}^{\star}(\nabla r_{1:t-k-1}(\bm{x}_{t-k})-\bm{g}_{t-k:t}),$
where $t-k$ denotes the most recent step at which we chose to prune. The last equality holds by the definition of $\bm{g}_{k}^{I}$ (assuming no predictions). Intuitively, we retain explicit gradient history only since the last pruning step $t-k$, while summarizing earlier history implicitly via the dual mapping of $\bm{x}_{t-k}$. The crux of the paper is showing that the way history is split (explicitly tracked after pruning, and implicitly captured before) is what controls dynamic regret.
@@ B.1 The Strong Dynamic Optimistic FTRL Lemma
$\bm{x}_{t+1}\doteq\operatorname*{argmin}_{\bm{x}}h_{0:t}(\bm{x})+\tilde{f}_{t+1}(\bm{x})$
@@
Since $\tilde{f}_{T+1}(\cdot)$ does not affect the algorithm, we can set it to 00. Rearranging:
@@ B.2 Bounding the First Part (I):
[…]
@@
[…]
It follows that $\bm{x}_{t}$ satisfies the optimality condition for $h_{0:t-1}(\bm{x})+\langle{\bm{\tilde{g}}_{t}},{\bm{x}}\rangle+\langle{\bm{g}^{I}_{t}},{\bm{x}}\rangle$ and hence is a minimizer [^10] Since $h_{0:t}(\cdot)$ is strongly convex, the minimizer of the two expressions in the lemma statement always exist and unique. In the case $\sigma_{1:t}=0$, we abuse notation by writing ``$=\operatorname*{argmin}"$ instead of ``$\in\operatorname*{argmin}"$. This is not problematic because we do not require the equivalence of the minimizers. thereof. ∎
@@
where $(a)$ follows by dropping the negative $-r_{t}(\bm{x}_{t+1})$ term and $(b)$ by defining $\bm{y}_{t}\doteq\arg\min_{\bm{x}}h_{0:t-1}(\bm{x})+\langle{\bm{p}_{t}},{\bm{x}}\rangle$. Then,
@@
$=h_{0:t-1}(\bm{x}_{t})+\langle{\bm{p}_{t}},{\bm{x}}\rangle-h_{0:t-1}(\bm{x}_{t+1})-\langle{\bm{p}_{t}},{\bm{x}_{t+1}}\rangle-r_{t+1}(\bm{x}_{t+1})$ (90)
@@
Where again the inequality follows by dropping the non-positive $r_{t+1}(\cdot)$. Next, we invoke Lemma D.1 with
@@
[…]
@@ B.3 Bounding the Second Part (II):
where we chose the 00 vector from $\mathcal{N}_{\mathcal{X}}(\bm{u}_{t+1})$. Hence, the length of $\bm{q}_{t}$ satisfies:
@@
Note that the negative term in (106) admits a tight lower bound of 00 and will therefore be omitted. Next, we derive a bound on the state vector’s length $\|\bm{p}_{1:t}\|$.
@@
(Optimistically Bounded State) Let $\{\bm{p}_{t}\}_{t=1}^{T}$ be a sequence of vectors such that each $\bm{p}_{t}=\bm{g}_{t}+\bm{g}_{t}^{I}$, where $\bm{g}_{t}\in\partial f_{t}(\bm{x}_{t})$, and $\bm{g}t^{I}\in\partial I_{\mathcal{X}}(\bm{x}_{t})$ is chosen according to the construction in (11). That is, $\bm{p}_{t}$ corresponds to the linearization of the composite function $f_{t}(\cdot)+I_{\mathcal{X}}(\cdot)$ around $\bm{x}_{t}$. Then, for any $t$, the following holds:
@@ Appendix C Missing Proofs for Section 3
The cumulative-squared prediction error: $E_{T}\doteq\sum_{t=1}^{T}\epsilon_{t}^{2},\quad\epsilon_{t}\doteq\|\bm{g}_{t}-\bm{\tilde{g}_{t}}\|.$ (109)
The path length: $P_{T}\doteq\sum_{t=1}^{T-1}\|\bm{u}_{t+1}-\bm{u}_{t}\|$ (110)
The prediction-weighted path length: $H_{T}\doteq\sum_{t=1}^{T-1}\epsilon_{t}\|\bm{u}_{t+1}-\bm{u}_{t}\|$ (111)
@@
where the last equality holds from the definition of $[t]^{+}$. That is, for the slot $\tau$ when the $\max$ evaluates to 00, we write it as $\left(\sqrt{\frac{E_{\tau}}{P^{\prime}_{\tau}}}-\sqrt{\frac{E_{\tau-1}}{P^{\prime}_{\tau-1}}}\right)+\left(\sqrt{\frac{E_{\tau-1}}{P^{\prime}_{\tau-1}}}-\sqrt{\frac{E_{\tau}}{P^{\prime}_{\tau}}}\right)$. Now, by telescoping
@@
[…]
@@ C.3.1 Bounding the $A_{T}$ term
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{t=1}^{T}\left[\|\bm{u}_{t+1}-\bm{u}_{t}\|\sqrt{E_{t}}\left(\sqrt{\frac{1}{P^{\prime}_{1}}}-\sqrt{\frac{1}{P^{\prime}_{t}}}\right)\right]$ (167)
=$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{t=1}^{T}\left[\|\bm{u}_{t+1}-\bm{u}_{t}\|\sqrt{E_{t}}\sqrt{\frac{1}{P^{\prime}_{1}}}\right]$ (168)
@@
where $(a)$ is from $E_{\tau}\geq E_{\tau-1}$ for all $\tau$ by definition; $(b)$ holds similarly because $E_{\tau}$ is non-decreasing on $\tau$ (or $t$); $(c)$ holds because $P^{\prime}_{\tau-1}\leq P^{\prime}_{\tau}$ for all $\tau$ (for any slot $t$) and hence we can add additional positive terms and create the entire sum from $\tau=2$ to $\tau=t$, instead of only the partial sum of terms in $[t]^{+}$; $(d)$ holds by writing the telescoping sum; $(e)$ holds by dropping the last term which is negative; and finally in $(f)$ we used the fact that $E_{T}\geq E_{t},\forall t$.
@@
Clarification on the term “recursive”: In previous regularization strategies, the strong convexity at time $t$, $\sigma_{1:t}$, can be expressed in closed form as $\sigma_{1:t}=\sigma\sqrt{E_{t}}$. This follows from defining each $\sigma_{t}$ to be exactly $\sigma(\sqrt{E_{t}}-\sqrt{E_{t-1}})\leq\frac{\sigma\epsilon_{t}^{2}}{2\sqrt{E_{t-1}}}$. However, when $\sigma_{t}$ is defined more generally as a scalar $\sigma\delta_{t}$, where $\delta_{t}$ can take any value in $[0,\sigma(\frac{\epsilon_{t}^{2}}{2\sqrt{E_{t-1}}})]$, we lose this compact form. Instead, $\sigma_{1:t}$ is now recursively defined as $\sigma_{1:t-1}+\sigma\delta_{t}$. As discussed, this ensures minimal regularization, impacting both the algorithm’s behavior and the analysis.
@@
where $(a)$ follows by Lemma 4.3 which bounds $\|\bm{p}_{1:t}\|$, and by the fact that $\sigma_{1:t-1}\leq\sigma_{1:t}$, $(b)$ by dropping the non-negative $-r_{t}(\bm{x}_{t+1})$, $(c)$ by the definition of $\delta_{t}$, $(d)$ by bounding each $\|\bm{u}_{t}\|$ in $r_{t}(\bm{u}_{t}),t\leq T$ by $R$, $(e)$ since $\sigma=\nicefrac{{1}}{{8R^{2}}}$.
@@ Appendix D Auxiliary Lemmas
[…]
@@
(Orabona, 2022, Section 7.6) Let $R,\sigma$ be positive real numbers that satisfy $\sigma\leq\frac{1}{2b}$, $b>0$. Then, for any $a_{t}\geq 0,\forall t\in[T]$:
$\sum_{t=1}^{T}\min\left(\frac{a_{t}^{2}}{2\sigma\sqrt{\sum_{\tau=1}^{t-1}a_{\tau}}},ba_{t}\right)\leq\frac{\sqrt{2}}{\sigma}\sqrt{\sum_{t=1}^{T}a_{t}}.$ (191)
@@
The proof of this lemma closely follows the argument presented in the cited section. However, due to the inclusion of the term $R$ in the second part of the $\min$, we adapt the proof accordingly to account for this difference.
$\sum_{t=1}^{T}\min\left(\frac{a_{t}^{2}}{2\sigma\sqrt{\sum_{\tau=1}^{t-1}a_{\tau}^{2}}},ba_{t}\right)\leq\sum_{t=1}^{T}\sqrt{\min\left(\frac{a_{t}^{4}}{4\sigma^{2}\sum_{\tau=1}^{t-1}a_{\tau}^{2}},b^{2}a_{t}^{2}\right)}$ (192)
@@ Appendix E Comparison with the Literature
  • (Scroccaro et al., 2023) Obtains a bound of $\mathcal{O}((P_{T}+1)\sqrt{D_{T}+1})$, where $D_{T}\doteq\|\sum_{t=1}^{T}\nabla f_{t}(\bm{x}_{t-1})-\nabla\tilde{f}_{t}(\bm{x}_{t-1})\|$. which is $\mathcal{O}(P_{T})$ in the best case.
@@
  • This work achieves a dynamic regret bound of $\mathcal{R}_{T}=\mathcal{O}((1+P_{T})\sqrt{E_{T}}$, which is 00 in the best case.
@@
  • (Scroccaro et al., 2023) Obtains a bound of $\mathcal{O}(\sqrt{(P_{T}+1)(D_{T}+\theta_{T})})$, where $\theta_{T}$ is the sum of corrective terms $\theta_{t}$, which are added to the learning rate to ensure monotonicity, and it holds that in the perfect gradient prediction case $\theta_{T}=\mathcal{O}(1+P_{T})$ and the regret bound is also $\mathcal{O}(1+P_{T})$.
@@
  • This work achieves a dynamic regret bound of $\mathcal{O}((1+\sqrt{P_{T}})\sqrt{E_{T}}+A_{T})$, which is 00 in the best case.

### v2

@@ A.4 Dual-perspective
$r^{\star}(\bm{y})\doteq\sup_{\bm{x}}\langle{\bm{x}},{\bm{y}}\rangle-r(\bm{x}).$ (62)
@@
$\bm{x}^{\text{RL}}_{t+1}=\arg\min_{x}\langle\bm{g}_{1:t},\bm{x}\rangle+r_{0:t}(\bm{x})=\nabla r_{0:t}^{\star}(-\bm{g}_{1:t}),$ (63)
@@
$\bm{x}_{t+1}=\arg\min_{x}\langle\bm{p}_{1:t},\bm{x}\rangle+r_{0:t}(\bm{x})=\nabla r_{0:t}^{\star}(-\bm{p}_{1:t})=\nabla r_{0:t}^{\star}(\nabla r_{1:t-k-1}(\bm{x}_{t-k})-\bm{g}_{t-k:t}),$ (64)
where tkt-k denotes the most recent step at which we chose to prune. The last equality holds by the definition of $\bm{g}_{k}^{I}$ (assuming no predictions). Intuitively, we retain explicit gradient history only since the last pruning step tkt-k, while summarizing earlier history implicitly via the dual mapping of $\bm{x}_{t-k}$. The crux of the paper is showing that the way history is split (explicitly tracked after pruning, and implicitly captured before) is what controls dynamic regret.
@@ B.1 The Strong Dynamic Optimistic FTRL Lemma
$\bm{x}_{t+1}\doteq\operatorname*{argmin}_{\bm{x}}h_{0:t}(\bm{x})+\tilde{f}_{t+1}(\bm{x})$ (66)
@@
Since $\tilde{f}_{T+1}(\cdot)$ does not affect the algorithm, we can set it to 0. Rearranging:
@@ B.2 Bounding the First Part (I):
We bound $\textbf{(I)}:h_{0:t}(\bm{x}_{t})-h_{0:t}(\bm{x}_{t+1})-r_{t}(\bm{x}_{t})$ in two ways, which results in the $\min(\cdot,\cdot)$ term. We present in this subsection Lemmas B.2 and 4.2, corresponding to each argument of the $\min(\cdot,\cdot)$.
@@
In case $(ii)$, Note that $(\bm{p}_{1:t-1}+\bm{\tilde{g}}_{t}+\sigma_{1:t-1}\bm{x}_{t})=-\nabla h_{1:t-1}(\bm{x}_{t})-\bm{\tilde{g}}_{t}$, and hence (81) reduces to $0\in\mathcal{N}_{\mathcal{X}}(\bm{x}_{t})$, which is always true. Thus, $\bm{x}_{t}$ satisfies (81) in this case too.
It follows that $\bm{x}_{t}$ satisfies the optimality condition for $h_{0:t-1}(\bm{x})+\langle{\bm{\tilde{g}}_{t}},{\bm{x}}\rangle+\langle{\bm{g}^{I}_{t}},{\bm{x}}\rangle$ and hence is a minimizer [^10] Since $h_{0:t}(\cdot)$ is strongly convex, the minimizer of the two expressions in the lemma statement always exist and unique. In the case $\sigma_{1:t}=0$, we abuse notation by writing $\operatorname*{argmin}$" instead of $\in\operatorname*{argmin}$". This is not problematic because we do not require the equivalence of the minimizers. thereof. ∎
@@
where $(a)$ follows by dropping the negative $r_{t}(\bm{x}_{t+1})$ term and $(b)$ by defining $\bm{y}_{t}\doteq\arg\min_{\bm{x}}h_{0:t-1}(\bm{x})+\langle{\bm{p}_{t}},{\bm{x}}\rangle$. Then,
@@
$=h_{0:t-1}(\bm{x}_{t})+\langle{\bm{p}_{t}},{\bm{x}}\rangle-h_{0:t-1}(\bm{x}_{t+1})-\langle{\bm{p}_{t}},{\bm{x}_{t+1}}\rangle-r_{t}(\bm{x}_{t+1})$ (94)
@@
Where again the inequality follows by dropping the non-positive $r_{t}(\cdot)$. Next, we invoke Lemma D.1 with
@@
$\textbf{(I)}\leq\min\left(\frac{1}{2}\|\bm{g}_{t}-\bm{\tilde{g}}_{t}\|_{t-1,*}^{2},2R\epsilon_{t}\right)$ (102)
@@ B.3 Bounding the Second Part (II):
where we chose the 0 vector from $\mathcal{N}_{\mathcal{X}}(\bm{u}_{t+1})$. Hence, the length of $\bm{q}_{t}$ satisfies:
@@
Note that the negative term in (110) admits a tight lower bound of 0 and will therefore be omitted. Next, we derive a bound on the state vector’s length $\|\bm{p}_{1:t}\|$.
@@
(Optimistically Bounded State) Let $\{\bm{p}_{t}\}_{t=1}^{T}$ be a sequence of vectors such that each $\bm{p}_{t}=\bm{g}_{t}+\bm{g}_{t}^{I}$, where $\bm{g}_{t}\in\partial f_{t}(\bm{x}_{t})$, and $\bm{g}_{t}^{I}\in\partial I_{\mathcal{X}}(\bm{x}_{t})$ is chosen according to the construction in (13). That is, $\bm{p}_{t}$ corresponds to the linearization of the composite function $f_{t}(\cdot)+I_{\mathcal{X}}(\cdot)$ around $\bm{x}_{t}$. Then, for any $t$, the following holds:
@@ Appendix C Missing Proofs for Section 3
The cumulative-squared prediction error: $E_{T}\doteq\sum_{t=1}^{T}\epsilon_{t}^{2},\quad\epsilon_{t}\doteq\|\bm{g}_{t}-\bm{\tilde{g}_{t}}\|.$
The path length: $P_{T}\doteq\sum_{t=1}^{T-1}\|\bm{u}_{t+1}-\bm{u}_{t}\|\ignorespaces\ignorespaces$
The prediction-weighted path length: $H_{T}\doteq\sum_{t=1}^{T-1}\epsilon_{t}\|\bm{u}_{t+1}-\bm{u}_{t}\|\ignorespaces\ignorespaces$
@@
where the last equality holds from the definition of $t]^{+}$. That is, for the slot $\tau$ when the $\max$ evaluates to 0, we write it as $\left(\sqrt{\frac{E_{\tau}}{P^{\prime}_{\tau}}}-\sqrt{\frac{E_{\tau-1}}{P^{\prime}_{\tau-1}}}\right)+\left(\sqrt{\frac{E_{\tau-1}}{P^{\prime}_{\tau-1}}}-\sqrt{\frac{E_{\tau}}{P^{\prime}_{\tau}}}\right)$. Now, by telescoping
@@
where $(a)$ follows by Lemma 4.3 which bounds $\|\bm{p}_{1:t}\|$, and the fact that $\sigma_{1:t-1}\leq\sigma_{1:t}$, $(b)$ by bounding each $\|\bm{u}_{t}\|$ in $r_{t}(\bm{u}_{t}),t\leq T$ by $R$, $(c)$ by using Lemma C.3 for the first sum, and by selecting [^11] Note that $\bm{u}_{T+1}$ does not affect the algorithm and hence we can set it without loss of generality $\bm{u}_{T+1}$ such that $\|\bm{u}_{T+1}-\bm{u}_{T}\|\geq\frac{R}{4}$, which allows us to append the $\frac{R^{2}}{2}\sigma_{1:T}$ term to the second sum as the summand with index $T$, $(d)$ by using Lemma C.4 to re-write $\sigma_{1:t}$, $(e)$ since $P^{\prime}_{t}$ is non-decreasing, $(f)$ from $P^{\prime}_{t}\geq 2R$, $(g)$ by $\sigma=\nicefrac{{1}}{{2\sqrt{2R}}}$ and Lemma D.3, $(h)$ by the fact that $2R\geq\|\bm{u}_{t+1}-\bm{u}_{t}\|,\forall t$, and finally $(i)$ by Lemma D.2 with $a_{t}=\|\bm{u}_{t+1}-\bm{u}_{t}\|$. ∎
@@ C.3.1 Bounding the $A_{T}$ term
$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{t=1}^{T}\left[\|\bm{u}_{t+1}-\bm{u}_{t}\|\sqrt{E_{t}}\left(\sqrt{\frac{1}{P^{\prime}_{1}}}-\sqrt{\frac{1}{P^{\prime}_{t}}}\right)\right]$ (168)
$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{t=1}^{T}\left[\|\bm{u}_{t+1}-\bm{u}_{t}\|\sqrt{E_{t}}\sqrt{\frac{1}{P^{\prime}_{1}}}\right]$ (169)
@@
where $(a)$ is from $E_{\tau}\geq E_{\tau-1}$ for all $\tau$ by definition; $(b)$ holds similarly because $E_{\tau}$ is non-decreasing on $\tau$ (or $t$); $(c)$ holds because $P^{\prime}_{\tau-1}\leq P^{\prime}_{\tau}$ for all $\tau$ (for any slot $t$) and hence we can add additional positive terms and create the entire sum from $\tau=2$ to $\tau=t$, instead of only the partial sum of terms in $t]^{+}$; $(d)$ holds by writing the telescoping sum; $(e)$ holds by dropping the last term which is negative; and finally in $(f)$ we used the fact that $E_{T}\geq E_{t},\forall t$.
@@
Clarification on the term “recursive”: In previous regularization strategies, the strong convexity at time $t$, $\sigma_{1:t}$, can be expressed in closed form as $\sigma_{1:t}=\sigma\sqrt{E_{t}}$. This follows from defining each $\sigma_{t}$ to be exactly $\sigma(\sqrt{E_{t}}-\sqrt{E_{t-1}})\leq\frac{\sigma\epsilon_{t}^{2}}{2\sqrt{E_{t-1}}}$. However, when $\sigma_{t}$ is defined more generally as a scalar $\sigma\delta_{t}$, where $\delta_{t}$ can take any value in $0,\sigma(\frac{\epsilon_{t}^{2}}{2\sqrt{E_{t-1}}})]$, we lose this compact form. Instead, $\sigma_{1:t}$ is now recursively defined as $\sigma_{1:t-1}+\sigma\delta_{t}$. As discussed, this ensures minimal regularization, impacting both the algorithm’s behavior and the analysis.
@@
where $(a)$ follows by Lemma 4.3 which bounds $\|\bm{p}_{1:t}\|$, and by the fact that $\sigma_{1:t-1}\leq\sigma_{1:t}$, $(b)$ by dropping the non-negative $r_{t}(\bm{x}_{t+1})$, $(c)$ by the definition of $\delta_{t}$, $(d)$ by bounding each $\|\bm{u}_{t}\|$ in $r_{t}(\bm{u}_{t}),t\leq T$ by $R$, $(e)$ since $\sigma=\nicefrac{{1}}{{8R^{2}}}$.
@@ Appendix D Auxiliary Lemmas
(McMahan, 2017, Lemma 7) Given a convex function $\phi_{1}(\cdot)$ with a minimizer $\bm{y}_{1}\doteq\operatorname*{argmin}_{\bm{y}}{\phi_{1}(\bm{y})}$ and a function $\phi_{2}(\cdot)=\phi_{1}(\cdot)+\psi(\cdot)$ that is $1$-strongly convex w.r.t some norm $\|\cdot\|$. Then,
@@
(Orabona, 2022, Section 7.6) Let $\sigma$ be a positive real number that satisfy $\sigma\leq\frac{1}{2b}$, $b>0$. Then, for any $a_{t}\geq 0,\forall t\in[T]$:
$\sum_{t=1}^{T}\min\left(\frac{a_{t}^{2}}{2\sigma\sqrt{\sum_{\tau=1}^{t-1}a_{\tau}}},ba_{t}\right)\leq\frac{\sqrt{2}}{\sigma}\sqrt{\sum_{t=1}^{T}a^{2}_{t}}.$ (192)
@@
The proof of this lemma closely follows the argument presented in the cited section. However, due to the inclusion of the term $b$ in the second part of the $\min$, we adapt the proof accordingly to account for this difference.
$\sum_{t=1}^{T}\min\left(\frac{a_{t}^{2}}{2\sigma\sqrt{\sum_{\tau=1}^{t-1}a_{\tau}^{2}}},ba_{t}\right)=\sum_{t=1}^{T}\sqrt{\min\left(\frac{a_{t}^{4}}{4\sigma^{2}\sum_{\tau=1}^{t-1}a_{\tau}^{2}},b^{2}a_{t}^{2}\right)}$ (193)
@@ Appendix E Comparison with the Literature
  • (Scroccaro et al., 2023) Obtains a bound of $\mathcal{O}((P_{T}+1)\sqrt{D_{T}+1})$, where $D_{T}\doteq\|\sum_{t=1}^{T}\nabla f_{t}(\bm{y}_{t-1})-\nabla\tilde{f}_{t}(\bm{y}_{t-1})\|$. which is $\mathcal{O}(P_{T})$ in the best case.
@@
  • This work achieves a dynamic regret bound of $\mathcal{R}_{T}=\mathcal{O}((1+P_{T})\sqrt{E_{T}}$, which is 0 in the best case.
@@
  • (Scroccaro et al., 2023) Obtains a bound of $\mathcal{O}(\sqrt{(P_{T}+1)(D_{T}+\theta_{T})})$, where $\theta_{T}$ is the sum of corrective terms $\theta_{t}$, which are added to the learning rate to ensure monotonicity, and it holds that in the perfect gradient prediction case $\theta_{T}=\mathcal{O}(1+P_{T})$ and the regret bound is also $\mathcal{O}(P_{T})$.
@@
  • This work achieves a dynamic regret bound of $\mathcal{O}((1+\sqrt{P_{T}})\sqrt{E_{T}}+A_{T})$, which is 0 in the best case.